I don't know what this type of equation is called but: 4x+3y=70
4x-27=80
How would you solve this?

Answer by Alan3354(69443)

4x+3y=70
4x-27=80 --> 4x = 107 --> x = 107/4
-----
4x+3y=70
4(107/4) + 3y = 70
107 + 3y = 70
3y = - 37
y = -37/3